In a certain Barangay, 550 random residents are chosen in which 225 are males. At a=0.01, test the hypothesis that more than 35% of the Barangay population are male residents.

1
Expert's answer
2022-04-29T05:41:43-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p\le0.35"

"H_a:p>0.35"

This corresponds to a right-tailed test, for which a z-test for one population proportion will be used.

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{\dfrac{225}{550}-0.35}{\sqrt{\dfrac{0.35(1-0.353)}{550}}}\approx2.9054"

The p-value is "p =P(Z>0.0970)= 0.001834"

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a right-tailed test is "z_c = 2.3263."

The rejection region for this right-tailed test is "R = \{z: z> 2.3263\}."

Since it is observed that "z = 2.9054 \ge2.3263= z_c ," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value is "p =0.922726," and since "p = 0.001834 \le 0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population proportion "p" is greater than 0.35, at the "\alpha = 0.01" significance level.